Suppose that you invest in a two-year Treasury bond with a coupon rate of 6% and $1,000 par. Suppose that you buy this bond at a price of exactly $1,000. You intend to hold this bond to maturity and reinvest the coupons until the bond matures. You expect to reinvest the coupons in an account that pays an APR of 2.01%, with semi-annual compounding. What is the effective annual rate of return on your investment? Hint: see Example 8 in the Lecture.

Do not round at intermediate steps in your calculation. Express your answer in percent. Round to three decimal places. Do not type the % symbol. If the return is negative, then include a minus sign.
Answer #1

SInce price of Bond is same as its par value,Hence coupon rate is same as YTM
YTM =6%
Coupon =6%*1000 =60
FV of Coupon at 2.01% reinvestment rate =60*(1+2.01%) =61.206
Amount received at of 2 years =FV of Coupon at 2.01% reinvestment rate+Coupon+Par Value
=61.206+60+1000 =1121.206
Effective annual acaRate =(FV/PV)^n-1 =(1121.206/1000)^(1/2)-1 =5.89%